Athanasios Dedes

University of Bonn, Germany

SUSY02, June 17-23, 2002, DESY Hamburg, Germany

B_{s} → μ^{+}μ^{−} and Trilepton Events

at the Tevatron

In collaboration with:

H. Dreiner and U. Nierste, Phys. Rev. Lett. 87, 251804 (2001),

                    hep-ph/0108037

and H. Dreiner, U. Nierste, and P. Richardson, to appear

1. The decay B_{s} → μ^{+}μ^{−} in the MSSM

2. B_{s} → μ^{+}μ^{−} and its correlation with the Trilepton events at Run II, (g − 2)_{μ} and Light Higgs Boson mass in the mSUGRA scenario.

3. Conclusions

See also SUSY02 talks by

J. Urban, R. Arnowitt, C. Kolda, C. Balazs, S. Baek, W. de Boer.

B_{s} → μ^{+}μ^{−} and Trilepton Events at the Tevatron  1

The Experimental Facts

√ LEP Higgs boson mass bound (preliminary)

M_{h} ≥ 107(113) GeV sin^{2}(β − α) ≃ 1  (1)

√ The muon anomalous magnetic moment

−4.2 × 10^{−10} ≤ δa_{μ}^{SUSY} ≤ 41.3 × 10^{−10} @90% CL (2)

√ The CDF upper bound on B_{s}-meson decay to dimuons

B(B_{s} → μ^{+}μ^{−}) < 2.0 × 10^{−6} @90% CL  (3)

√ B(b → sγ) [!conservative bound!]

2.0 × 10^{−4} ≤ B(b → sγ) ≤ 5.0 × 10^{−4}  (4)

√ SUSY searches

...... , m_{}_{χ̃_{1}^{±}} > ∼ 100 GeV, ........  (5)

B_{s} → μ^{+}μ^{−} and Trilepton Events at the Tevatron  2

B_{s} → μ^{+}μ^{−} vs. δa_{μ}^{SUSY} in the MSSM

• δa_{μ}^{SUSY} depends upon 7 parameters (T. Moroi)

δa_{μ}^{SUSY} = f(M_{1} , M_{2} , A_{μ} , μ , tan β , mμ̃_{L} , mμ̃_{R})

• B(B_{s} → μ^{+}μ^{−}) depends (mainly) upon 7 parameters

B(B_{s} → μ^{+}μ^{−}) = f(M_{2} , A_{t} , μ , tan β , M_{A} , mt̃_{L} , mt̃_{R})

• Approximately :

B(B_{s} → μ^{+}μ^{−}) ≃ 3 × 10^{−7}[\frac{m_{μ}M_{B_{s}}}{M_{W}^{2}}]^{2} \frac{tan^{6} β}{M_{A}^{4}}μ^{2}A_{t}^{2} C(mt̃, mχ̃) ,

MSSM : K. S. Babu and C. Kolda, ’99, C. Huang, W. Liao and Q. Yan, ’99 P. H. Chankowski and L. Slawianowska, ’01, C. Bobeth, T. Ewerth, F. Kruger and J. Urban, ’01, ’02 G. Isidori, A. Retico, ’01, C. Bobeth, A. J. Buras, F. Kruger and J. Urban, Dec. ’01,

mSUGRA : A. Dedes, H. K. Dreiner and U. Nierste, ’01 R. Arnowitt, B. Dutta, T. Kamon and M. Tanaka, ’02 H. Baer, C. Balazs, A. Belyaev, J. K. Mizukoshi, X. Tata and Y. Wang,’02

GMSB, AMSB : S. w. Baek, P. Ko and W. Y. Song, ’02

• The Maximum mass for the Lightest Observable SUSY par-

ticle (LOSP) [Feng and Matchev ’01]

δa_{μ}^{SUSY} ≃ 18 × 10^{−10} \frac{tan β}{50} (\frac{603 GeV}{M_{LOSP}^{max}})^{2} ,

• The Maximum mass for the Heaviest Higgs Boson

B(B_{s} → μ^{+}μ^{−}) ≃ 5 × 10^{−7} (\frac{tan β}{50})^{6} (\frac{550 GeV}{M_{A}^{max}})^{4}

University of Bonn  3  Athanasios Dedes

B_{s} → μ^{+}μ^{−} and Trilepton Events at the Tevatron  4

B_{s} → μ^{+}μ^{−} vs. (g − 2)_{μ} in mSUGRA

A. Dedes, H. K. Dreiner and U. Nierste, ’01

Solid : B(B_{s} → μ^{+}μ^{−}) Dashed : δa_{μ}^{SUSY} in 10^{−10} units Dot-dashed : Higgs mass Shaded : Excluded

B_{s} → μ^{+}μ^{−} and Trilepton Events at the Tevatron  5

A. Dedes, H. K. Dreiner and U. Nierste, ’01

Solid : B(B_{s} → μ^{+}μ^{−}) Dashed : δa_{μ}^{SUSY} in 10^{−10} units Dot-dashed : Higgs mass Shaded : Excluded

B_{s} → μ^{+}μ^{−} and Trilepton Events at the Tevatron  6

Trileptons and B_{s} → μ^{+}μ^{−} at Run II

• The “Gold-plated” Mode • Final State : lll, l^{±}l^{±}, l^{±}l^{∓}τ, l^{±}ττ, τττ, τ^{±}τ^{±} l ≡ e, μ , focus on 3L ≡ lll √ New and improved analysis compared to the one by Matchev and Pierce ’99 → slightly bigger signal over the Background

B_{s} → μ^{+}μ^{−} and Trilepton Events at the Tevatron  8

A. Dedes, H. Dreiner, U. Nierste, P. Richardson ’02

Tevatron 5σ reach for 3L and L = 2fb^{−1} Tevatron 5σ reach for 3L and L = 10fb^{−1} Tevatron 5σ reach for 3L and L = 30fb^{−1} Dashed : δa_{μ}^{SUSY} in 10^{−10} units Dotted : Light Higgs Boson mass Tevatron 5σ reach for 3L and L = 2fb^{−1} Tevatron 5σ reach for 3L and L = 10fb^{−1} Tevatron 5σ reach for 3L and L = 30fb^{−1} Solid : B(B_{s} → μ^{+}μ^{−}) Dashed : δa_{μ}^{SUSY} in 10^{−10} units Dotted : Light Higgs Boson mass Tevatron 5σ reach for 3L and L = 2fb^{−1} Tevatron 5σ reach for 3L and L = 10fb^{−1} Tevatron 5σ reach for 3L and L = 30fb^{−1} Solid : B(B_{s} → μ^{+}μ^{−}) Dashed : δa_{μ}^{SUSY} in 10^{−10} units Dotted : Light Higgs Boson mass

Conclusions

√ If B_{s} → μ^{+}μ^{−} is observed at the Tevatron with L = 2fb^{−1} [B(B_{s} → μ^{+}μ^{−}) > ∼ 5 × 10^{−7}] then :

\underline{MSSM}

• the heaviest Higgs boson mass is bounded from above in the MSSM, M_{A} < ∼ 550 GeV

• upper bound on the ratio A_{t}/M_{A}

• strong support for the minimal SUSY SO(10) model

\underline{mSUGRA}

√ If B_{s} → μ^{+}μ^{−} is observed at the Tevatron Run II [B(B_{s} → μ^{+}μ^{−}) > ∼ 10^{−8}] then :

• δa_{μ}^{SUSY} deviates significantly from zero

• m_{h}^{max} < ∼ 120 GeV

• B_{s} → μ^{+}μ^{−} complements the trilepton searches

• a No-Lose scenario for mSUGRA at the Tevatron with M_{1/2} , M_{0} < ∼ 400 and tan β > ∼ 40

It might be that these are the SUSY footprints
at Tevatron.......